OpenStudy (anonymous):

Can you show me how to calculate: b^2b^1/3 / b^1/2

OpenStudy (turingtest):

remember \[x^ax^b=x^{a+b}\]and\[x^a/x^b=x^{a-b}\]then just take care with the fractions: \[b^2b^{1/3}/b^{1/2}=b^{2+1/3-1/2}=b^{12/6+2/6-3/6}=b^{11/6}\]